OpenStudy (anonymous):

5x − 3(x − 8) = 4(x + 9) X=

OpenStudy (anonymous):

OK, like before, the first step is to multiply out and collect like terms. What do you get when you multiply -3(x-8)?

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

GoLearnMath that's really not useful

OpenStudy (anonymous):

-3x-21

OpenStudy (lilg132):

you expand the brackets then rearrange so you will get 5x − 3(x − 8) = 4(x + 9) expand it and you will get 5x - 3x + 24 = 4x + 36 bring the x's to one side and numbers on the other 5x - 3x - 4x = 36 - 24 -2x = 12 -x = 6 x = -6

OpenStudy (lilg132):

dont forget to click good answer :)

OpenStudy (anonymous):

close! Remember that when you multiply two negative numbers by each other, the negatives cancel. try again?

OpenStudy (saifoo.khan):

clicked. ;)

OpenStudy (lilg132):

ktklown are you talking about me?

OpenStudy (anonymous):

No I'm trying to teach math.

OpenStudy (anonymous):

To the original asker.

OpenStudy (anonymous):

and lilg your answer was wrong for the last question ilovethis21 has asked

OpenStudy (lilg132):

yes thanks for pointing that out

OpenStudy (anonymous):

ilovethis21 -- I'd like to help you work through this if you'd like. let's try again to expand... -3(x-8)?

OpenStudy (anonymous):

GoLearnMath actuall you got it wrong, sorry